# LLM-Mediated Guidance of MARL Systems

### v1

@@ Abstract
In complex multi-agent environments, achieving efficient learning and desirable behaviours is a significant challenge for Multi-Agent Reinforcement Learning (MARL) systems. This work explores the potential of combining MARL with Large Language Model (LLM)-mediated interventions to guide agents toward more desirable behaviours. Specifically, we investigate how LLMs can be used to interpret and facilitate interventions that shape the learning trajectories of multiple agents. We experimented with two types of interventions, referred to as controllers: a Natural Language (NL) Controller and a Rule-Based (RB) Controller. The NL Controller, which uses an LLM to simulate human-like interventions, showed a stronger impact than the RB Controller. Our findings indicate that agents particularly benefit from early interventions, leading to more efficient training and higher performance. Both intervention types outperform the baseline without interventions, highlighting the potential of LLM-mediated guidance to accelerate training and enhance MARL performance in challenging environments.
@@ 1 Introduction
Cooperative MARL research has developed techniques to effectively optimize collective return in simulated environments (Rashid et al., 2020; Yuan et al., 2023; Albrecht et al., 2024). This enables the deployment of multi-agent systems (MAS) that can efficiently solve complex tasks, particularly in tasks that factorize into parallel subtasks and/or take place in the physical world (e.g., robotics) and can benefit from spatially-scattered agents (Calvaresi et al., 2021). However, what if the reward function is misspecified? This can happen because the reward is difficult to define in a way that avoids reward hacking (Skalse et al., 2022). Alternatively, what if the test time environment or system goals change slightly? We would like a user to be able to steer a MARL system towards more desirable behaviour (human-in-the-loop). These are all key challenges that arise in real-world domains. In addition, we do not want to assume the user is a MARL expert. Ideally, the user could steer the system in an intuitive and simple way. Therefore, we consider steering a MAS using natural language. The user issues high-level strategies that an LLM then translates into actions to communicate with the MAS. While examples of humans intervening and controlling static programs/interfaces via LLMs are pervasive (Hong et al., 2023), we know of fewer examples controlling single-agent learning systems and no examples controlling MA learning systems.
@@
Integrating LLMs with RL presents exciting opportunities for enhancing agent performance, particularly in complex MA environments. Instruction-aligned models with advanced reasoning and planning capabilities are well-suited for this task. Prompted correctly, these models provide real-time, context-aware strategies, guiding agents through challenges where traditional RL methods struggle, especially in environments with large action/observation spaces or sparse rewards, particularly during early training. We envision a future where LLM-RL combinations can manage increasingly dynamic environments, with LLMs handling complex interactions and dynamically changing observation and action spaces. Our research explores this potential in MARL. We allow users to quickly ’fine-tune’ a base MARL system by guiding the agents using free-form natural language or rule-based interventions in the training process. This adaptation helps the system align more closely with the user’s bespoke task requirements, ensuring that agents develop behaviours tailored to the challenges of the environment. We have specifically chosen the Aerial Wildfire Suppression (AWS) environment from the HIVEX suite (Siedler, 2025)  [^1] Environment: https://github.com/hivex-research/hivex-environments
@@
The AWS environment presents dynamic and high-stakes cooperative scenarios, where the unpredictability of wildfire spread creates an evolving challenge. Factors such as wind direction, humidity, terrain slope, and temperature—hidden from the agents—add layers of complexity. Solving this environment requires seamless collaboration among agents, where strategic coordination is essential to containing fires. With AWS, users engage in a problem simulating real-world wildfire management. The combination of a physically and visually rich simulation, open-ended scenarios and environmental conditions makes AWS a demanding environment and a great challenge.
In this work, we test whether combining current MARL and LLM techniques can allow users to steer and guide a MARL system towards more desirable behaviour in the challenging AWS environment. We consider two users: the simple Rule-Based (RB) Controller and a more sophisticated Natural Language (NL) Controller. The NL Controller simulates how humans might interact with the MAS, i.e., in free-form natural language. We compare these against our baseline, a setup with no test-time interventions. We summarize our core contributions as follows:
@@ 2 Related Work
Integrating LLMs into RL has become pivotal for enhancing agent performance in complex environments. Advanced LLMs, specifically, their instruction fine-tuned versions, have demonstrated significant capabilities in providing high-level guidance, common-sense reasoning, and strategic planning, thereby possibly improving RL agents’ adaptability and generalization (Bubeck et al., 2023). Recent works, such as those by Wang et al. (2023) and Chiang & Lee (2023), have shown that LLMs can assist RL agents by mediating natural language instructions and guiding behaviours, especially in environments where traditional reward signals are sparse or ineffective (Kajić et al., 2020). However, these studies primarily focus on single-agent scenarios or environments with relatively straightforward dynamics. In contrast, our work emphasizes MA environments with complex, interdependent dynamics, demonstrating that LLM-driven interventions can significantly accelerate learning in such settings.
@@
Historically, human-in-the-loop RL involved human feedback in guiding the learning process (Kamalaruban et al., 2019). LLMs have emerged as scalable, real-time alternatives, providing domain-specific knowledge and policy suggestions to correct suboptimal behaviours (Chiang & Lee, 2023). While previous research by Narvekar et al. (2020) explored dynamic curriculum approaches, where models generate instructions that change based on the agent’s progress, our approach leverages LLMs not for curriculum generation but for real-time human and LLM-based interventions specifically designed to address the challenges of coordinating multiple agents. This key distinction significantly impacts the effectiveness of the learning process in more complex environments. LLMs also address challenges in long-term planning and common-sense reasoning (Hao et al., 2023) by offering early and intermediate guidance that traditional RL methods often lack. Previous studies in robotics have similarly leveraged LLMs as high-level strategic planners, enabling more effective decision-making in tasks that require long-term coordination and planning (Tang et al., 2023; Ahn et al., 2022). While these works illustrate the potential of LLMs in improving decision-making in tasks requiring extended sequences of actions, our work expands this concept by integrating LLM-driven interventions at critical points in the learning process, specifically in MA scenarios where coordinated action over long horizons is crucial.
@@
In MA systems, LLMs show promise in improving coordination and strategic planning. Traditional MARL approaches, like MADDPG and QMIX, face limitations due to the complexity of joint action spaces and sparse rewards (Lowe et al., 2017; Rashid et al., 2018). Other work specifies a mediator to steer an MA system towards a desirable equilibrium without incorporating any LLM (Zhang et al., 2024). While recent works, such as Kwon et al. (2023), have demonstrated that a global reward can control an MA system with a single intervention at the beginning—showing how to cheaply design a reward model in natural language using an LLM—these approaches do not fully address the dynamic nature of MA environments where frequent adaptations are necessary (Wang et al., 2024). Our research builds on these insights by demonstrating that periodic LLM interventions significantly enhance cooperation and learning efficiency, especially in dynamic and unpredictable environments such as AWS. This adaptive intervention strategy addresses the shortcomings of static coordination approaches by providing real-time guidance that aligns with the evolving state of the environment and agent interactions.
@@
LLM interventions offer adaptive guidance that complements traditional policy shaping (Griffith et al., 2013), evolving with the learning process. Our method does not fit neatly into Open-loop or Closed-loop categories (Sun et al., 2024), as it temporarily replaces RL agent actions with LLM-guided interventions in both NL and RB setups. Unlike prior work using LLMs for agent communication and collaboration, our approach uniquely employs a central LLM to craft high-level strategies for coordinating multiple agents. This aligns with open research directions, specifically ”language-enabled Human-in/on-the-Loop Frameworks” (Sun et al., 2024), by mimicking human-in-the-loop strategies. In contrast to Wang et al. (2023), which focuses on building agent capabilities, we emphasize centralized LLM-driven strategy development. Whether through strategic foresight or moment-to-moment decision-making, our approach adapts to dynamic environments. Assuming we only compare the inference cost of our LLM-Mediator module, we gain an advantage as long as its cost is lower than the total inference cost of the agent over deployment.
@@ 3 The Aerial Wildfire Suppression Environment
The AWS environment presents a rich and challenging scenario for AI agents, far exceeding the simplicity of traditional grid-based worlds. Unlike grid worlds, which offer limited spatial complexity, this environment presents a three-dimensional, continuous, and dynamic landscape where agents must adapt to fire spread patterns that are difficult to predict. AWS is built in Unity (Juliani et al., 2020), a game development engine, offering a saturated, semi-realistic-looking visual component compared to Atari-like environments (Mnih et al., 2013), providing a more complex and high-dimensional observation space with both feature vector and visual data. This diversity of input, combined with the need for real-time decision-making and collaboration, makes it a robust and challenging platform for testing advanced AI strategies in complex, non-deterministic scenarios.
The AWS environment simulates a complex scenario where agents must manage and mitigate the spread of wildfires. This environment is designed to challenge agents with complex decision-making tasks, requiring both individual action and coordinated teamwork. The main focus is on reducing fire spread, protecting key assets, the village, and navigating a large, bounded terrain. The agent’s primary objective is to minimize the fire’s burning duration by extinguishing as many burning trees as possible and preparing unburned areas to prevent further spread. Agents can either extinguish burning trees or redirect the fire’s path by preparing/wetting the surrounding forest area.
The environment includes three agents, each with a feature vector ($\mathbb{R}^{8}$) and visual-observation space ($42\times 42$ RGB grid). Feature vector observations include agent 2-d position, direction, a binary indicator of whether the agent is holding water, position of the nearest tree, and the nearest tree’s state, burning or not burning. The agents move at a constant velocity with actions to steer left, right, and drop water if held. They operate within a bounded area on an island. A negative reward is given if the agent crosses the environment’s boundary. Water surrounds the island; steering the aeroplane toward and collecting it produces a positive reward. Agents earn positive rewards for extinguishing or preparing forest areas to slow fire spread and for extinguishing the wildfire completely. Detailed environment specifications A.4, detailed task list, reward breakdown and calculations can be found in the Appendix in Reward Description and Calculation A.5.
@@ 4 Intervention Controllers and LLM-Mediator
Our system supports interventions from two types of controllers: the Rule-Based (RB) and Natural Language (NL) Controller, which differ in their level of sophistication for generating interventions. The RB Controller uses predefined rules and a prompt template, producing rudimentary agent instructions. In contrast, the NL Controller communicates in free-form natural language, mimicking human behaviour. This allows it to generate more complex strategies and contextually relevant guidance. The LLM-Mediator processes both types of interventions, translating them and temporarily overwriting the agents’ learned policy actions, guiding them to complete specific tasks (Figure 5). This framework enables adaptive guidance and control in dynamic environments (Figure 4).
@@ 4.2 Natural Language Controller
The NL Controller uses a prompt template with partial feature vector observation data (Figure 6). This information is provided as a list of all agents’ observations and descriptions in natural language (Figure 7). The observation information formatted prompt is provided to an LLM, mimicking human behaviour, which generates a strategy directing agents to specific map locations. The NL Controller’s high level directive is to “Develop a strategy to extinguish all fires”. The resulting strategy is then passed to the LLM-Mediator. Matching with the Rule-Based Controller, the LLM-Mediator processes this more sophisticated strategy and returns agent-readable actions.
@@ 4.3 Mediator
At the core, controllers act as prompt crafters. When a controller intervention prompt is issued, it is sent to the LLM-Mediator. Once the LLM-Interpreter processes the intervention, a task list is generated for each agent, and a 300-time-step cooldown period begins. During this period, agents are assigned their first task, and actions are generated to guide them toward task completion. These actions overwrite the agents’ policy actions, such as steering left or right. If the agent holds water during the intervention period, the LLM-Mediator ensures it is retained by default. Each task includes a key to identify the agent and specify a target location (Figure 8). As long as the target location is not reached, actions continue to be auto-generated and issued to the agent. If the task is not completed within 300 time steps, a new intervention can be triggered. Figure 6 illustrates a basic terminal interface, as we imagine a human controller or firefighter using it to review observations, in combination with camera feed and radar data, etc., to determine whether an intervention should be issued.
@@ 4.4 Pseudocode: MARL with LLM Interventions
Input: Multi-agent environment, PPO policy $\pi_{\theta}$, LLM-Mediator, intervention frequency $f$
Initialize environment, cooldown timers $\{c^{i}\leftarrow f\}_{i=1}^{N}$ for all agents $i$ and policy parameters $\theta_{0}$
for $episode=1,2,\ldots$ do
     Reset environment and cooldown timers $\{c^{i}\leftarrow f\}_{i=1}^{N}$
     while not done do
         Collect observations $\{s_{t}^{i}\}_{i=1}^{N}$ for all agents $i\in\{1,\ldots,N\}$
         Compute actions $\{a_{t}^{i}\}_{i=1}^{N}$ using policy $\pi_{\theta}(s_{t}^{i})$ for each agent $i$
         for each agent $i\in\{1,\ldots,N\}$ do
              if $c^{i}==f$ then
                  Generate intervention using LLM-Mediator:
                      $a_{t}^{i}\leftarrow\text{LLM-Mediator}(s_{t}^{i})$
                  Reset cooldown timer for agent $i$: $c^{i}\leftarrow f$
              else if agent is currently following an LLM task then
                  Decrement cooldown timer: $c^{i}\leftarrow c^{i}-1$
                  if $c^{i}<0$ then
                       Reset cooldown timer: $c^{i}\leftarrow f$
@@
         Perform a single step in the environment:
             $\{s_{t+1}^{i},r_{t}^{i}\}_{i=1}^{N}\leftarrow\text{env.step}(\{a_{t}^{i}\}_{i=1}^{N})$
         Store transitions $\{(s_{t}^{i},a_{t}^{i},r_{t}^{i},s_{t+1}^{i})\}_{i=1}^{N}$ for all agents
@@
     Update PPO policy $\pi_{\theta}$:
         Combine transitions from all agents into a shared buffer
         Compute advantage estimates $\{\hat{A_{t}^{i}}\}_{i=1}^{N}$ and rewards-to-go $\{\hat{R_{t}^{i}}\}_{i=1}^{N}$
         Optimize PPO objective to get $\theta_{k+1}$
@@
Algorithm 1 Multi-Agent RL with LLM Interventions and Cooldown Timers

### v2

@@ Abstract
In complex multi-agent environments, achieving efficient learning and effective coordination remains a major challenge for Multi-Agent Reinforcement Learning (MARL) systems. This work explores the integration of MARL with Large Language Model (LLM)-mediated guidance to steer agents toward improved collective behaviour. We investigate a centralized intervention framework in which an LLM temporarily overrides agent actions during training to provide sparse, high-level guidance. We evaluate two types of controllers: a Rule-Based (RB) Controller and a Natural Language (NL) Controller that uses a small (7B/8B) instruction-tuned LLM to simulate human-like guidance. Experiments in the Aerial Wildfire Suppression (AWS) environment show that both controllers outperform a no-intervention baseline, with RB interventions yielding the strongest gains. Notably, interventions applied during early training stages significantly accelerate learning and lead to higher final performance. Our results demonstrate that LLM-mediated interventions can improve coordination and training efficiency in challenging MARL settings without modifying reward functions or policies, highlighting their potential as a practical mechanism for human-inspired steering of multi-agent systems.
MARL, LLMs, Reinforcement Learning, Multi-Agent Reinforcement Learning
@@ 1 Introduction
Cooperative MARL research has developed techniques to effectively optimize collective return in simulated environments (Rashid et al., 2020; yuan2023survey; albrecht2024multi). This enables the deployment of multi-agent systems (MAS) that can efficiently solve complex problems, particularly those that factorize into parallel subtasks and/or occur in the physical world (e.g., robotics), where spatially distributed agents are beneficial (Calvaresi et al., 2021). However, what if the reward function is misspecified? This can happen because the reward is difficult to define in a way that avoids reward hacking (Skalse et al., 2022). Alternatively, what if the test time environment or system goals change slightly? We would like a user to be able to steer a MARL system towards more desirable behaviour (human-in-the-loop). While our experiments use an LLM to simulate human interventions, the Natural Language Controller interface is designed to be directly usable by human operators, allowing seamless replacement of the LLM with human intervention. These are all key challenges that arise in real-world domains. In addition, we do not want to assume the user is a MARL expert. Ideally, the user could steer the system in an intuitive and simple way. Therefore, we consider steering a MAS using natural language. The user issues high-level strategies that an LLM then translates into actions to communicate with the MAS. While examples of humans intervening and controlling static programs/interfaces via LLMs are pervasive (Hong et al., 2023), we know of fewer examples controlling single-agent learning systems and no examples controlling MA learning systems.
@@
Integrating LLMs with RL presents exciting opportunities for enhancing agent performance, particularly in complex MA environments. Instruction-aligned models with advanced reasoning and planning capabilities are well-suited for this task. When prompted correctly, they provide real-time, context-aware strategies that guide agents through challenges where traditional RL methods struggle. This is especially valuable in environments with large action and observation spaces or sparse rewards, particularly during early training. We envision a future where LLM-RL combinations can manage increasingly dynamic environments, with LLMs handling complex interactions and dynamically changing observation and action spaces. Our research explores this potential in MARL. We allow users to quickly training a MARL system by guiding the agents using free-form natural language or rule-based interventions in the training process. This adaptation helps the system align more closely with the user’s bespoke task requirements, ensuring that agents develop behaviours tailored to the challenges of the environment. We have specifically chosen the Aerial Wildfire Suppression (AWS) environment from the HIVEX suite (siedler_hivex_2025)  [^1] Environment: https://github.com/hivex-research/hivex-environments
@@
The AWS environment presents a dynamic and high-stakes cooperative scenario, where the unpredictability of wildfire spread creates an evolving challenge. Factors such as wind direction, humidity, terrain slope, and temperature – hidden from the agents – add layers of complexity. Solving this environment requires seamless collaboration among agents, where strategic coordination is essential to containing fires. With AWS, users engage in a problem simulating real-world wildfire management. The combination of a physically and visually rich simulation, open-ended scenarios and environmental conditions makes AWS a demanding environment and a great challenge.
In this work, we test whether combining MARL and LLM techniques can allow users to steer and guide a MARL system towards more desirable behaviour in the challenging AWS environment. We consider two users: the simple Rule-Based (RB) Controller and a more sophisticated Natural Language (NL) Controller. The NL Controller simulates how humans might interact with the MAS, i.e., in free-form natural language. We compare these against our baseline, a setup with no interventions. Unlike prior work that uses LLMs for reward design, curriculum generation, or agent communication, our method introduces a centralized LLM-mediated intervention mechanism that temporarily overrides learned MARL policies without updating them directly, enabling sparse, human-like steering during training without modifying rewards or policies. We summarize our core contributions as follows:
@@ 2 Related Work
Integrating LLMs into RL has become pivotal for enhancing agent performance in complex environments. Advanced LLMs, specifically, their instruction fine-tuned versions, have demonstrated significant capabilities in providing high-level guidance, common-sense reasoning, and strategic planning, thereby possibly improving RL agents’ adaptability and generalization (Bubeck et al., 2023). Recent works, such as those by Wang et al. (2023) and Chiang and Lee (2023), have shown that LLMs can assist RL agents by mediating natural language instructions and guiding behaviours, especially in environments where traditional reward signals are sparse or ineffective (Kajić et al., 2020). However, these studies primarily focus on single-agent scenarios or environments with relatively straightforward dynamics. In contrast, our work emphasizes MA environments with complex, interdependent dynamics, demonstrating that LLM-driven interventions can significantly accelerate learning in such settings.
@@
Historically, human-in-the-loop RL has relied on human feedback to guide the learning process (Kamalaruban et al., 2019). LLMs have emerged as scalable, real-time alternatives, providing domain-specific knowledge and policy suggestions to correct suboptimal behaviours (chiang_lee_2023). While prior work explores dynamic curriculum approaches, where models generate instructions based on agent progress (Narvekar et al., 2020), our method instead uses LLMs for real-time interventions designed specifically to support coordination in multi-agent systems. This key distinction significantly impacts the effectiveness of the learning process in more complex environments. LLMs also address challenges in long-term planning and common-sense reasoning (Hao et al., 2023) by offering early and intermediate guidance that traditional RL methods often lack. Previous studies in robotics have similarly leveraged LLMs as high-level strategic planners, enabling more effective decision-making in tasks that require long-term coordination and planning (Tang et al., 2023; Ahn et al., 2022). While these works illustrate the potential of LLMs in improving decision-making in tasks requiring extended sequences of actions, our work expands this concept by integrating LLM-driven interventions at critical points in the learning process, specifically in MA scenarios where coordinated action over long horizons is crucial.
@@
In MA systems, LLMs show promise in improving coordination and strategic planning. Traditional MARL approaches, like MADDPG and QMIX, face limitations due to the complexity of joint action spaces and sparse rewards (Lowe et al., 2017; Rashid et al., 2018). Other work specifies a mediator to steer an MA system towards a desirable equilibrium without incorporating any LLM (Zhang et al., 2024). While recent works, such as Kwon et al. (2023), have demonstrated that a global reward can control an MA system with a single intervention at the beginning – showing how to cheaply design a reward model in natural language using an LLM – these approaches do not fully address the dynamic nature of MA environments where frequent adaptations are necessary (Wang et al., 2024). Our research builds on these insights by demonstrating that periodic LLM interventions significantly enhance cooperation and learning efficiency, especially in dynamic and unpredictable environments such as AWS. This adaptive intervention strategy addresses the shortcomings of static coordination approaches by providing real-time guidance that aligns with the evolving state of the environment and agent interactions.
@@
LLM interventions offer adaptive guidance that complements traditional policy shaping (Griffith et al., 2013), evolving with the learning process. Our method does not fit neatly into open-loop or closed-loop categories (Sun et al., 2024), as it temporarily replaces RL agent actions with LLM-guided interventions in both NL and RB setups. Unlike prior work using LLMs primarily for agent communication or capability building, our approach employs a centralized LLM to generate high-level strategies for coordinating multiple agents. This aligns with emerging research directions in language-enabled human-in/on-the-loop frameworks (Sun et al., 2024), while focusing on direct intervention rather than agent-side reasoning.
While prior work leverages LLMs for curriculum generation (Sun et al., 2024), reward design via language (Kwon et al., 2023), or improving agent autonomy and exploration (Wang et al., 2023), these approaches do not intervene directly at the action level of trained MARL policies. In contrast, our method introduces a centralized LLM-mediated intervention mechanism that temporarily overrides learned agent actions during training, without modifying reward functions or policy parameters, enabling sparse, human-like steering in complex multi-agent environments.
@@ 3 The Aerial Wildfire Suppression Environment
The Aerial Wildfire Suppression (AWS) environment presents a rich and challenging testbed for AI agents, substantially exceeding the simplicity of traditional grid-based worlds. Unlike grid environments with limited spatial structure, AWS features a continuous, three-dimensional, and dynamic landscape in which agents must adapt to stochastic and often unpredictable fire-spread dynamics. Built in Unity (Juliani et al., 2020), the environment offers a semi-realistic visual presentation that contrasts with Atari-style benchmarks (Mnih et al., 2013), resulting in a high-dimensional observation space composed of both vector-based features and raw visual inputs. This combination of visual complexity, real-time decision-making, and multi-agent coordination makes AWS a demanding platform for evaluating advanced AI methods in non-deterministic, safety-critical scenarios.
The environment simulates a large, bounded island on which agents are tasked with managing and mitigating an active wildfire. The primary objective is to minimize the total burning duration by extinguishing burning trees and proactively preparing surrounding forest areas to slow or redirect fire spread. Achieving this goal requires a combination of individual decision-making and coordinated teamwork across agents, as they must balance fire suppression, navigation, and resource collection while protecting key assets such as a nearby village.
AWS consists of three agents, each controlling an aerial firefighting aeroplane. Agents receive a feature vector observation space ($\mathbb{R}^{8}$) and a visual observation space consisting of a $42\times 42$ RGB image captured by a downward-facing camera. Vector observations include the agent’s two-dimensional position and direction, an indicator of whether the agent is carrying water, the location of the nearest tree, and the burning state of that tree. Agents move at a constant velocity and can steer continuously while executing a discrete action to drop water when available. The operational area is bounded; crossing the boundary incurs a negative reward.
Water surrounds the island and can be collected by navigating toward it, yielding a positive reward. Additional positive rewards are granted for extinguishing burning trees, wetting unburned trees to reduce future fire spread, and fully suppressing the wildfire. Together, these reward signals encourage agents to balance immediate suppression with longer-term containment strategies. Detailed environment specifications are provided in Section A.2, and a full task list and reward definitions in the Appendix Section A.3.
@@ 4 Intervention Controllers and LLM-Mediator
Our system supports interventions from two types of controllers: the Rule-Based (RB) and Natural Language (NL) Controller, which differ in their level of sophistication for generating interventions. The RB Controller uses predefined rules and a prompt template, producing rudimentary agent instructions. In contrast, the NL Controller communicates in free-form natural language, mimicking human behaviour. This allows it to generate more complex strategies and contextually relevant guidance. The LLM-Mediator processes both types of interventions, translating them and temporarily overwriting the agents’ learned policy actions, guiding them to complete specific tasks (Figure 5). This framework enables adaptive guidance and control in dynamic environments (Figure 4). Importantly, LLM-issued actions are not treated as expert demonstrations, but rather as regular transitions, and policy updates remain fully on-policy PPO updates.
@@ 4.2 Natural Language Controller
The NL Controller uses a prompt template with partial feature vector observation data (Figure 6). This information is provided as a list of all agents’ observations and descriptions in natural language (Figure 7). The observation information formatted prompt is provided to an LLM, mimicking human behaviour, which generates a strategy directing agents to specific map locations. The NL Controller’s high level directive is to “Develop a strategy to extinguish all fires”. The resulting strategy is then passed to the LLM-Mediator. Matching with the Rule-Based Controller, the LLM-Mediator processes this more sophisticated strategy and returns agent-readable actions. Although we use an LLM to generate natural language strategies in our experiments, the NL Controller mirrors a human-facing interface and can be directly operated by a human issuing free-form instructions.
@@ 4.3 Mediator
At the core, controllers act as prompt crafters. When a controller intervention prompt is issued, it is sent to the LLM-Mediator. Once the LLM-Mediator processes the intervention, a task list is generated for each agent, and a 200-time-step cooldown period begins. We empirically found this cooldown length to balance intervention effectiveness and policy autonomy; shorter cooldowns led to unstable behavior due to frequent overrides, while longer cooldowns reduced responsiveness to changing environmental conditions. During this period, agents are assigned their first task, and actions are generated to guide them toward task completion. These actions overwrite the agents’ policy actions, such as steering left or right. If the agent holds water during the intervention period, the LLM-Mediator ensures it is retained by default. Each task includes a key to identify the agent and specify a target location (Figure 8). As long as the target location is not reached, actions continue to be auto-generated and issued to the agent. If the task is not completed within 200 time steps, a new intervention can be triggered. Figure 6 illustrates a basic terminal interface, as we imagine a human controller or firefighter using it to review observations, in combination with camera feed and radar data, etc., to determine whether an intervention should be issued.
@@ 4.4 Pseudocode: MARL with LLM Interventions
Algorithm 1 Multi-Agent PPO with LLM interventions and cooldown scheduling
  Environment env; PPO policy $\pi_{\theta}$; LLM mediator $M$; intervention frequency $f$; agents $i=1..N$
  Initialize policy parameters $\theta\leftarrow\theta_{0}$
  for each training episode do
  Reset environment
  Initialize cooldown timers $c^{i}\leftarrow f$ for all agents
  while episode not finished do
   Observe states $s_{t}^{i}$ and sample actions $a_{t}^{i}\sim\pi_{\theta}(\cdot\mid s_{t}^{i})$
   for each agent $i$ do
    if cooldown expired ($c^{i}=f$) then
     Override action with LLM guidance: $a_{t}^{i}\leftarrow M(s_{t}^{i})$
     Reset cooldown timer $c^{i}\leftarrow f$
    else if agent is executing an LLM-issued task then
     Decrease cooldown timer $c^{i}\leftarrow c^{i}-1$
     if cooldown underflows then
     Reset cooldown timer $c^{i}\leftarrow f$
@@
   Step environment with actions $\{a_{t}^{i}\}_{i=1}^{N}$ and observe rewards and next states
   Store transitions $(s_{t}^{i},a_{t}^{i},r_{t}^{i},s_{t+1}^{i})$
@@
  Update PPO policy using shared experience buffer